# DGC-GNN: Descriptor-free Geometric-Color Graph Neural Network for 2D-3D Matching

### v1

@@ Abstract
Direct matching of 2D keypoints in an input image to a 3D point cloud of the scene without requiring visual descriptors has garnered increased interest due to its lower memory requirements, inherent privacy preservation, and reduced need for expensive 3D model maintenance compared to visual descriptor-based methods. However, existing algorithms often compromise on performance, resulting in a significant deterioration compared to their descriptor-based counterparts. In this paper, we introduce DGC-GNN, a novel algorithm that employs a global-to-local Graph Neural Network (GNN) that progressively exploits geometric and color cues to represent keypoints, thereby improving matching robustness. Our global-to-local procedure encodes both Euclidean and angular relations at a coarse level, forming the geometric embedding to guide the local point matching. We evaluate DGC-GNN on both indoor and outdoor datasets, demonstrating that it not only doubles the accuracy of the state-of-the-art descriptor-free algorithm but, also, substantially narrows the performance gap between descriptor-based and descriptor-free methods. The code and trained models will be made publicly available.
@@ 1 Introduction
Establishing 2D-3D matches plays a crucial role in various computer vision applications, including visual localization irschara2009structure ; sattler2012improving ; sattler2018benchmarking ; sarlin2019coarse ; wang2021continual ; sarlin2020superglue , 3D reconstruction snavely2006photo ; cui2015global ; schoenberger2016sfm ; lindenberger2021pixel , and Simultaneous Localization and Mapping (SLAM) eade2006scalable ; mur2015orb ; mur2017orb . Traditional methods for establishing point-to-point matches involve extracting keypoints and descriptors from a query image, then matching the 2D and 3D descriptors using exhaustive search. To circumvent the computationally expensive matching process, some approaches irschara2009structure ; sarlin2019coarse narrow the search space by employing image retrieval methods nister2006scalable ; Arandjelovic16 first to identify the most similar images in the database, and then perform descriptor-based image matching lowe2004distinctive ; detone2018superpoint ; dusmanu2019d2 ; sarlin2020superglue ; shi2022clustergnn between the query and retrieved images. The 2D-3D correspondences are subsequently established by connecting the 2D-2D image matches with the prebuilt 2D-3D correspondences in the database. Another approach sattler2011fast proposes an efficient vocabulary-based method to directly build 2D-to-3D matches by searching through all point descriptors. Sattler et al. sattler2012improving ; sattler2016efficient further explore the combination of both 2D-3D and 3D-2D search as an active correspondence search step for a faster and more efficient matching process.
While descriptor-based algorithms achieve state-of-the-art accuracy, they store and maintain high-dimensional visual descriptors for each point in potentially large 3D point clouds, significantly increasing both memory footprint and maintenance costs. The stored model often requires orders of magnitude more storage than the point cloud and images alone zhou2022geometry . These methods are susceptible to privacy attacks dosovitskiy2016inverting ; dosovitskiy2016generating ; chelani2021privacy and necessitate computationally expensive model maintenance and descriptor update procedures zhou2022geometry when incorporating new descriptors or points into the model. Several approaches have been proposed to address these limitations. Yang et al.Yang_2022_CVPR employ a learned point selection module to sample a subset of the point cloud for scene compression. Other methods li2020hierarchical ; brachmann2021dsacstar directly learn a function that maps 2D pixels to 3D coordinates without explicitly storing the 3D scene. Additionally, ng2022ninjadesc introduces an adversarial learning framework to develop content-concealing visual descriptors that prevent privacy leakage.
Recently, researchers campbell2020solving ; liu2020learning have begun exploring deep learning techniques for cross-domain direct 2D-3D matching and pose estimation without visual descriptors, showcasing the potential of descriptor-free matching through differentiable geometric optimization. The recently proposed GoMatch zhou2022geometry represents significant progress in descriptor-free 2D-3D matching, achieving reasonable matching performance on a variety of real-world datasets li2018megadepth ; shotton2013scene ; kendall2015posenet . GoMatch first identifies keypoints in the query image, which, along with the 3D points from the model, are converted to bearing vectors in the camera coordinate system. The algorithm employs the attention mechanism sarlin2020superglue ; vaswani2017attention to effectively establish reliable 2D-3D correspondences. While GoMatch attains reasonable accuracy, its performance still significantly lags behind its descriptor-based counterparts sarlin2020superglue ; sarlin2019coarse ; sattler2016efficient . Additionally, our experiments reveal that its reliance on geometric cues from the points and their local neighbors renders GoMatch incapable of distinguishing geometrically indistinct structures.
@@
Figure 1: 2D-3D matching (shown by green lines) with the proposed DGC-GNN and GoMatch zhou2022geometry . In this example, DGC-GNN obtains 78 correct matches with 0.02 meters camera translation and ${}^{\circ}$ rotation errors, while GoMatch finds only 17 inliers with a pose error of 0.37 meters and ${}^{\circ}$.
@@
Building upon these observations and the groundwork set by GoMatch, we propose a novel graph-based pipeline, named DGC-GNN, which leverages geometric and color cues in a global-to-local manner for descriptor-free 2D-3D matching. Our pipeline encodes both position and RGB information for each point and extracts a global distance-angular embedding to guide local point matching. Taking inspiration from shi2022clustergnn , we employ a cluster-based transformer to constrain information flow within local clusters. We observe, from real-world datasets, that the proposed DGC-GNN leads to substantial improvements in the number of correct matches and the accuracy of pose estimation. Notably, it doubles the accuracy of GoMatch, thereby reducing the gap between descriptor-based and descriptor-free methods. In summary, our paper makes the following contributions:
  • We introduce a descriptor-free global-to-local GNN for direct 2D-3D point matching. The network leverages multiple cues and incorporates a progressive clustering module to represent the keypoints. This pipeline enhances the robustness and accuracy of 2D-3D matching while requiring low memory, being privacy-preserving and free from 3D model maintenance.
  • We demonstrate that color information is crucial for 2D-3D matching. By incorporating RGB encoding into our network, we observe significant improvements in performance.
  • Extensive experiments on multiple real-world datasets show that the proposed network outperforms previous methods by a large margin on both matching and visual localization tasks.
2 Descriptor-Free 2D-3D Matching
@@ 2.1 Problem Formulation and Notation
Considering keypoints $\mathbf{P}=\{\mathbf{p}_{n}\in\mathbb{R}^{2}\;|\;n=1,...,N\}$ from a query image $I$ and a database 3D point cloud $\mathbf{Q}=\{\mathbf{q}_{m}\in\mathbb{R}^{3}\;|\;m=1,...,M\}$, where, optionally, each 3D point is associated with a visual descriptor $\mathbf{d}\in\mathbb{R}^{D}$. The task of 2D-3D point matching is to find a set $\mathcal{M}_{\mathbf{p,q}}$ of corresponding keypoints such that
@@
Bearing Vector. Similar to zhou2022geometry , we adopt bearing vectors as keypoint representation for both the 2D and 3D points to alleviate their cross-domain nature and represent them in the same space. The bearing vector is the direction from the camera center to a 3D point in the camera coordinate system. Given an image, a 2D pixel $\mathbf{p}_{n}$ is uplifted to bearing vector as $[\mathbf{b}_{\mathbf{p},n},1]^{\text{T}}=\mathbf{K^{-1}}[\mathbf{p}_{n},1]^{\text{T}},\mathbf{b}_{\mathbf{p},,}\in\mathbb{R}^{2}$, where $\mathbf{K}$ is the intrinsic camera matrix. Given a 3D point $\mathbf{q}_{m}$, the corresponding bearing vector is $[\mathbf{b}_{\mathbf{p},m},1]^{\text{T}}=\frac{\mathbf{Rq}_{m}+\mathbf{t}}{[\mathbf{Rq}_{m}+\mathbf{t}]_{z}}$, where $\mathbf{R}$ is the camera rotation and $\mathbf{t}$ is its translation in the world coordinate system and subscript $z$ denotes the third component of the 3D vector.
@@ 2.2 Network Architecture
Figure 2: Pipeline overview. For keypoints from the 2D image and 3D points from the point cloud, the proposed DGC-GNN (1) considers the bearing vectors and the color at each bearing vector as input. (2) It extracts the point-wise position and color features with two separate encoders and mixes the features as $\mathbf{f_{p}}$ and $\mathbf{f_{p}}$. (3) The bearing vectors are clustered into $K$ groups, and geometric graphs are built upon the clusters to extract the global-level geometric embeddings $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ and $\mathbf{\hat{f}}_{\mathbf{q}}^{gg}$. (4) We then concatenate $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ with $\mathbf{f_{p}}$ and $\mathbf{\hat{f}}_{\mathbf{q}}^{gg}$ with $\mathbf{f_{p}}$, and build a local graph at each point as self-attention. A cluster-based attention module is adopted to enhance the local features by forcing the message passing only with the most related features. A differentiable layer matches and optimizes the improved features to obtain score matrix $\mathcal{S}$. Finally, an outlier filtering network is applied to prune the matches with low confidence, leading to the final 2D-3D correspondences $\mathcal{M}_{final}$.
The proposed DGC-GNN applies a hierarchical mechanism to leverage both color and geometric cues in a global-to-local fashion. The overall pipeline is illustrated in  Fig.2. We initially employ two local feature extractors to encode RGB and position information for each point simultaneously (Sec. 2.2.1). Additionally, we cluster the points based on their distances and generate global graphs to obtain the global-level geometric embeddings (Sec. 2.2.2). Next, we concatenate the local point features with their corresponding global features and input them into the cluster-based local matching module to identify the initial matches (Sec. 2.2.3). Finally, we incorporate a classification network to filter out matches with low confidence to refine the initial matches (Sec.2.3).
@@ 2.2.1 Local Feature Extraction
To extract points-wise features from both the 2D keypoint set $\mathbf{P}$ and the 3D point cloud $\mathbf{Q}$, we consider the inputs as bearing vectors equipped with color information: $\mathcal{P}=\{\mathbf{b_{p},c_{p}}\}$ and $\mathcal{Q}=\{\mathbf{b_{q},c_{q}}\}$. Two ResNet-style point encoders he2016deep ; campbell2020solving , denoted as $\mathcal{F}_{b}$ and $\mathcal{F}_{c}$, are applied to extract position and color embeddings separately. We then obtain the local point features, $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$, as follows:
@@
The resulting point-wise features $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$ are vectors with dimensions $\mathbb{R}^{N\times d}$ and $\mathbb{R}^{M\times d}$ respectively, where $N$ and $M$ represent the number of keypoints in $\mathbf{P}$ and $\mathbf{Q}$, and $d$ denotes the dimensionality of the encoded features, e.g., $d=128$ for SIFT features lowe2004distinctive .
@@ 2.2.2 Global Geometric Guidance
Global context guidance has demonstrated its effectiveness in various computer vision tasks sun2021loftr ; li2020hierarchical ; yu2021cofinet ; qin2022geometric . Global context helps to differentiate local descriptors from similar structures or patches, thereby reducing ambiguity. However, most existing methods sun2021loftr ; qin2022geometric consider the outputs from different encoding layers as global and local features. This approach is not suitable for our scenario, as our input is sparse points. Downsampling the sparse point cloud results in losing distinctive geometric structures. Hence, we adopt cluster-based geometric encoding to extract global embeddings. As shown in Fig. 3.(a) and (c), the input bearing vectors, both in the image and in the point cloud, are first clustered into $K$ groups. The groups represent distinct clusters, each associated with a cluster center as the global position, denoted by $\mathbf{\hat{b}}_{\mathbf{p},k}\in\mathbb{R}^{2},k=1,...K$. The corresponding global embedding is obtained by computing the average of the point embeddings within the cluster, thus $\mathbf{\hat{f}}_{\mathbf{p},k}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{f}_{\mathbf{p},k},\ \mathbf{\hat{f}}_{\mathbf{p},k}\in\mathbb{R}^{d}$. The same is conducted on the 3D points to obtain $\mathbf{\hat{b}}_{\mathbf{q}}$ and $\mathbf{\hat{f}}_{\mathbf{q}}$.
@@
Global Geometric Graph. To aggregate and extract the geometric relations among the clusters, we propose a novel graph neural network that encodes both distance and angular cues; the basic GNN structure is built upon huang2021predator ; zhou2022geometry . In the following, we describe the graph building for the 2D global points set $\mathcal{\hat{P}}=\{\mathbf{\hat{b}_{p}},\mathbf{\hat{f}_{p}}\}$ and the same goes for $\mathcal{\hat{Q}}=\{\mathbf{\hat{b}_{q}},\mathbf{\hat{f}_{q}}\}$. Each cluster center point $\mathbf{\hat{b}}_{\mathbf{p},x}$ is connected to its $k$-NN neighbours ($k\leq K$) in the coordinate space, and $\xi_{\mathbf{p},(x,y)}$ is the edge between center point $\mathbf{\hat{b}}_{\mathbf{p},x}$ and $\mathbf{\hat{b}}_{\mathbf{p},y}$. We update the feature $\mathbf{\hat{f}}_{\mathbf{p},x}$ using the following equation:
${}^{(t+1)}\mathbf{\hat{f}}_{\mathbf{p},x}=\max_{\xi_{\mathbf{p},(x,y)}}\mathcal{H}_{g}(^{(t)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus(^{(t)}\mathbf{\hat{f}}_{\mathbf{p},x}-^{(t)}\mathbf{\hat{f}}_{\mathbf{p},y})),$ (3)
where the $\oplus$ denotes concatenation and $\mathcal{H}_{g}(*)$ is the linear projection with instance normalization ulyanov2016instance and a LeakyReLU function maas2013rectifier . The global feature $\mathbf{\hat{f}}_{\mathbf{p},x}$ is updated twice, and calculated as
$\mathbf{\hat{f}}_{\mathbf{p},x}^{g}=\mathcal{H}_{g}(^{(0)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus^{(1)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus^{(2)}\mathbf{\hat{f}}_{\mathbf{p},x}).$ (4)
Besides the distance embedding, inspired by qin2022geometric , we also adopt the angular embedding to obtain rotation-invariant geometric cues for the global point representation. To do so, we define the embedding on cluster triplets as shown in Fig. 3.(b). Given bearing vector $\mathbf{\hat{b}}_{\mathbf{p},z}$ and two of its neighbors $\mathbf{\hat{b}}_{\mathbf{p},x}$ and $\mathbf{\hat{b}}_{\mathbf{p},y}$, the angular embedding of $\mathbf{\hat{b}}_{\mathbf{p},z}$ is defined as follows:
@@
where $\operatorname{sine}(\cdot)$ is a sinusoidal function and $\sigma_{a}$ is a controller constant. We update the global geometric embedding $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ as an angular-aware attention mechanism as follows:
@@
$\operatorname{Att}(\mathbf{\hat{f}}_{\mathbf{p}}^{g},\mathbf{A}_{\mathbf{p}})=(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{V}}).\frac{(\mathbf{A}_{\mathbf{p}}\mathbf{W^{A}})(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{Q}})^{\text{T}}+(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{Q}})(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{K}})^{\text{T}}}{\sqrt{dim}}.$ (7)
$\mathbf{W^{A}},\mathbf{W^{Q}},\mathbf{W^{K}},\mathbf{W^{V}}\in\mathbb{R}^{d\times d}$ are the projection matrices of each item. The local point representation with global geometric embedding is as $\mathcal{\widetilde{P}}=\{\mathbf{b}_{\mathbf{p}},\mathbf{\widetilde{f}}_{\mathbf{p}}\}$ and $\mathbf{\widetilde{f}}_{\mathbf{p}}=\mathbf{f}_{\mathbf{p}}\oplus\mathbf{\hat{f}}_{\mathbf{p}}^{gg},\mathbf{\widetilde{f}}_{\mathbf{p}}\in\mathbb{R}^{N\times 2d}$. The operation for the bearing vectors obtained from point cloud $\mathcal{\hat{Q}}$ is the same and we get $\mathcal{\widetilde{Q}}=\{\mathbf{b}_{\mathbf{q}},\mathbf{\widetilde{f}}_{\mathbf{q}}\}$.

### v2

# DGC-GNN: Leveraging Geometry and Color Cues for Visual Descriptor-Free 2D-3D Matching

@@ Abstract
Matching 2D keypoints in an image to a sparse 3D point cloud of the scene without requiring visual descriptors has garnered increased interest due to its low memory requirements, inherent privacy preservation, and reduced need for expensive 3D model maintenance compared to visual descriptor-based methods. However, existing algorithms often compromise on performance, resulting in a significant deterioration compared to their descriptor-based counterparts. In this paper, we introduce DGC-GNN, a novel algorithm that employs a global-to-local Graph Neural Network (GNN) that progressively exploits geometric and color cues to represent keypoints, thereby improving matching accuracy. Our procedure encodes both Euclidean and angular relations at a coarse level, forming the geometric embedding to guide the point matching. We evaluate DGC-GNN on both indoor and outdoor datasets, demonstrating that it not only doubles the accuracy of the state-of-the-art visual descriptor-free algorithm but also substantially narrows the performance gap between descriptor-based and descriptor-free methods. [^1] The code and trained models are available at: https://github.com/AaltoVision/DGC-GNN-release.
@@ 1 Introduction
Establishing 2D-3D matches plays a crucial role in various computer vision applications, including visual localization [20, 40, 42, 37, 53, 38], 3D reconstruction [48, 8, 43, 25], and Simultaneous Localization and Mapping (SLAM) [15, 31, 30]. Traditional methods for establishing point-to-point matches involve extracting keypoints and descriptors from a query image, then matching the 2D and 3D descriptors using exhaustive search. To circumvent the computationally expensive matching process, some approaches [20, 37] narrow the search space by employing image retrieval methods [33, 1] first to identify the most similar images in the database, and then perform descriptor-based image matching [27, 11, 14, 38, 45] between the query and retrieved images. The 2D-3D correspondences are subsequently established by connecting the 2D-2D image matches with the prebuilt 2D-3D correspondences in the database. Another approach [39] is to build 2D-to-3D matches by searching through all point descriptors with an efficient vocabulary-based method. Sattler et al. [40, 41] further explore the combination of both 2D-3D and 3D-2D search as an active correspondence search step for a faster and more efficient matching process.
@@
Figure 1: 2D-3D matching (shown by green lines) with the proposed DGC-GNN and GoMatch [56]. In this example, DGC-GNN obtains 78 correct matches with 0.02 meters camera translation and ${}^{\circ}$ rotation errors, while GoMatch finds only 17 inliers with a pose error of 0.37 meters and ${}^{\circ}$.
While descriptor-based algorithms achieve state-of-the-art accuracy, they store and maintain high-dimensional visual descriptors for each point in potentially large 3D point clouds. The stored model often requires orders of magnitude more storage than the point cloud and images alone [56]. These methods are susceptible to privacy attacks [13, 12, 7, 34] and necessitate computationally expensive model maintenance and descriptor update procedures [56] when incorporating new descriptors or points into the model. Several approaches have been proposed to address these limitations. Yang et al. [54] employ learned point selection to sample a subset of the point cloud for scene compression. Other methods [23, 3] directly learn a function that maps 2D pixels to 3D coordinates without explicitly storing the 3D scene. Additionally, [32] introduces an adversarial learning framework to develop content-concealing descriptors that prevent privacy leakage.
Recently, researchers [5, 26] have begun exploring deep learning techniques for cross-domain direct 2D-3D matching and pose estimation without visual descriptors, showcasing the potential of descriptor-free matching through differentiable geometric optimization. The recently proposed GoMatch [56] represents significant progress in descriptor-free 2D-3D keypoints matching, achieving reasonable matching performance on a variety of real-world datasets [24, 46, 22]. GoMatch first identifies keypoints in the query image, which, along with the 3D points from the model, are converted to bearing vectors in the camera coordinate system. The algorithm employs an attention mechanism [38, 52] to establish reliable 2D-3D correspondences effectively. While GoMatch attains reasonable accuracy, its performance still significantly lags behind its descriptor-based counterparts [38, 37, 41]. Additionally, it relies on geometric cues only from the points and their local neighbors, rendering it incapable of distinguishing geometrically similar structures.
@@
Building upon these observations and the groundwork set by GoMatch, we propose a novel graph-based pipeline, named DGC-GNN, which leverages geometric and color cues in a global-to-local manner for descriptor-free 2D-3D matching. DGC-GNN encodes position and RGB information for each point and extracts a global distance-angular embedding to guide local point matching. Taking inspiration from [45], we employ a cluster-based transformer to constrain information flow within local clusters. We observe, from real-world datasets, that DGC-GNN leads to substantial improvements in the number of correct matches and the accuracy of pose estimation. Notably, it doubles the accuracy of GoMatch, thereby reducing the gap between descriptor-based and descriptor-free methods. In summary, our paper makes the following contributions:
  • We introduce a visual descriptor-free global-to-local GNN for direct 2D-3D keypoint matching. The network leverages multiple cues and incorporates a progressive clustering module to represent the keypoints. This pipeline enhances the accuracy of sparse 2D-3D matching while requiring low memory, being privacy-preserving, and low cost from 3D model maintenance.
  • We demonstrate that color information for each point is crucial for 2D-3D matching. By incorporating RGB encoding into our network, we observe significant performance improvements.
  • Extensive experiments on real-world datasets show that DGC-GNN outperforms previous methods by a large margin on both matching and visual localization tasks.
Figure 2: Pipeline overview. For keypoints from the 2D image and 3D points from the point cloud, the proposed DGC-GNN (1) considers the bearing vectors and the color at each bearing vector as input. (2) It extracts the point-wise position and color features with two separate encoders and mixes the features as $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$. (3) The bearing vectors are clustered into $K$ groups, and geometric graphs are built upon the clusters to extract the global-level geometric embeddings $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ and $\mathbf{\hat{f}}_{\mathbf{q}}^{gg}$. (4) We then concatenate $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ with $\mathbf{f_{p}}$ and $\mathbf{\hat{f}}_{\mathbf{q}}^{gg}$ with $\mathbf{f_{p}}$, and build a local graph at each point as self-attention. A cluster-based attention module is adopted to enhance the local features by forcing the message to pass only with the most related features. A differentiable layer matches and optimizes the improved features to obtain score matrix $\mathcal{S}$. Finally, an outlier rejection network is applied to prune the matches with low confidence, leading to the final 2D-3D correspondences $\mathcal{M}_{final}$.
2 Visual Descriptor-Free 2D-3D Matching
@@ 2.1 Problem Formulation and Notation
Given keypoints $\mathbf{P}=\{\mathbf{p}_{n}\in\mathbb{R}^{2}\;|\;n=1,...,N\}$ from query image $I$ and database 3D point cloud $\mathbf{Q}=\{\mathbf{q}_{m}\in\mathbb{R}^{3}\;|\;m=1,...,M\}$, where, optionally, each 3D point is associated with a visual descriptor $\mathbf{d}\in\mathbb{R}^{D}$. The task is to find a set $\mathcal{M}_{\mathbf{p,q}}$ of corresponding keypoints such that
@@
Bearing Vector. Similar to [56], we adopt bearing vectors as keypoint representation for both the 2D and 3D points to alleviate their cross-domain nature and represent them in the same space. The bearing vector is the direction from the camera center to a 3D point in the camera coordinate system. Given an image, a 2D pixel $\mathbf{p}_{n}$ is uplifted to bearing vector as $[\mathbf{b}_{\mathbf{p},n},1]^{\text{T}}=\mathbf{K^{-1}}[\mathbf{p}_{n},1]^{\text{T}},\mathbf{b}_{\mathbf{p},n}\in\mathbb{R}^{2}$, where $\mathbf{K}$ is the intrinsic camera matrix. Given a 3D point $\mathbf{q}_{m}$, the corresponding bearing vector is
$[\mathbf{b}_{\mathbf{q},m},1]^{\text{T}}=\frac{\mathbf{Rq}_{m}+\mathbf{t}}{[\mathbf{Rq}_{m}+\mathbf{t}]_{z}},$ (2)
where $\mathbf{R}$ is the camera rotation and $\mathbf{t}$ is its translation in the world coordinate system and subscript $z$ denotes the third component of the 3D vector.
@@ 2.2 Network Architecture
The proposed DGC-GNN applies a hierarchical mechanism to leverage color and geometric cues in a global-to-local fashion. The overall pipeline is illustrated in Fig.2. We initially employ two local feature extractors to encode RGB and position information for each point simultaneously (Sec. 2.2.1). Additionally, we cluster the points based on their distances and generate global graphs to obtain the global-level geometric embeddings (Sec. 2.2.2). Next, we concatenate the local point features with their corresponding global features and input them into the cluster-based local matching module to identify the initial matches (Sec. 2.2.3). Finally, we incorporate a classification network to filter out matches with low confidence to refine the initial matches (Sec.2.3).
@@ 2.2.1 Local Feature Extraction
To extract points-wise features from both the 2D keypoint set $\mathbf{P}$ and the 3D point cloud $\mathbf{Q}$, we consider the inputs as bearing vectors equipped with color information: $\mathcal{P}=\{\mathbf{b_{p},c_{p}}\}$ and $\mathcal{Q}=\{\mathbf{b_{q},c_{q}}\}$. Two ResNet-style point encoders [18, 5], denoted as $\mathcal{F}_{b}$ and $\mathcal{F}_{c}$, are applied to extract position and color embeddings separately. We then obtain the local point features, $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$, as follows:
@@
The resulting point-wise features $\mathbf{f_{p}}$ and $\mathbf{f_{q}}$ are vectors with dimensions $\mathbb{R}^{N\times d}$ and $\mathbb{R}^{M\times d}$ respectively, where $N$ and $M$ represent the number of keypoints in $\mathbf{P}$ and $\mathbf{Q}$, and $d$ denotes the dimensionality of the encoded features, e.g., $d=128$
@@ 2.2.2 Global Geometric Guidance
Global context guidance has demonstrated its effectiveness in various computer vision tasks [49, 23, 55, 36]. Global context helps to differentiate local descriptors from similar structures or patches, thereby reducing ambiguity. However, most existing methods [49, 36] consider the outputs from different encoding layers as global and local features. This approach is not suitable for our scenario, as our input is sparse points. Downsampling the sparse point cloud results in losing distinctive geometric structures. Hence, we adopt cluster-based geometric encoding to extract global embeddings. As shown in Fig. 3.(a) and (c), the input bearing vectors, both in the image and in the point cloud, are first clustered into $X$ groups. The groups represent distinct clusters, each associated with a cluster center as the global position, denoted by $\mathbf{\hat{b}}_{\mathbf{p},x}\in\mathbb{R}^{2},x=1,...X$. The corresponding global embedding is obtained as the average of the point embeddings within a cluster as $\mathbf{\hat{f}}_{\mathbf{p},x}=\frac{1}{P^{\prime}}\sum_{p^{\prime}=1}^{P^{\prime}}\mathbf{f}_{\mathbf{p},p^{\prime}}$, where $\mathbf{\hat{f}}_{\mathbf{p},x}\in\mathbb{R}^{d}$, $P^{\prime}$ is the number of points in the $x$. The same is conducted on the 3D points to obtain $\mathbf{\hat{b}}_{\mathbf{q}}$ and $\mathbf{\hat{f}}_{\mathbf{q}}$.
@@
Global Geometric Graph. To aggregate and extract the geometric relations among the clusters, we propose a novel graph neural network that encodes both distance and angular cues; the basic GNN structure is built upon [19, 56]. In the following, we describe the graph construction for the 2D global points set $\mathcal{\hat{P}}=\{\mathbf{\hat{b}_{p}},\mathbf{\hat{f}_{p}}\}$ and the same goes for $\mathcal{\hat{Q}}=\{\mathbf{\hat{b}_{q}},\mathbf{\hat{f}_{q}}\}$. Each cluster center point $\mathbf{\hat{b}}_{\mathbf{p},x}$ is connected to its $k$-NN neighbours ($k\leq K$) in the coordinate space, and $\xi_{\mathbf{p},(x,y)}$ is the edge between center points $\mathbf{\hat{b}}_{\mathbf{p},x}$ and $\mathbf{\hat{b}}_{\mathbf{p},y}$. We update the feature $\mathbf{\hat{f}}_{\mathbf{p},x}$ using the following equation:
${}^{(t+1)}\mathbf{\hat{f}}_{\mathbf{p},x}=\max_{\xi_{\mathbf{p},(x,y)}}\mathcal{H}_{g1}(^{(t)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus(^{(t)}\mathbf{\hat{f}}_{\mathbf{p},x}-^{(t)}\mathbf{\hat{f}}_{\mathbf{p},y})),$ (4)
where the $\oplus$ denotes concatenation and $\mathcal{H}_{g1}(*)$ is the linear projection with instance normalization [51] and a LeakyReLU function [28]. The $max$ operator applies to the $k$-NN neighbors. The global feature $\mathbf{\hat{f}}_{\mathbf{p},x}$ is updated twice, and calculated as
$\mathbf{\hat{f}}_{\mathbf{p},x}^{g}=\mathcal{H}_{g2}(^{(0)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus^{(1)}\mathbf{\hat{f}}_{\mathbf{p},x}\oplus^{(2)}\mathbf{\hat{f}}_{\mathbf{p},x}).$ (5)
$\mathcal{H}_{g2}$ has a similar structure to $\mathcal{H}_{g1}$, but without shared weights. Besides the distance embedding, inspired by [36], we also adopt the angular embedding to obtain rotation-invariant geometric cues for the global representation. To do so, we define the embedding on cluster triplets as shown in Fig. 3. (b). Given bearing vector $\mathbf{\hat{b}}_{\mathbf{p},x}$ and two of its neighbors $\mathbf{\hat{b}}_{\mathbf{p},y}$ and $\mathbf{\hat{b}}_{\mathbf{p},z}$, the angular embedding of <$\mathbf{\hat{b}}_{\mathbf{p},x}$, $\mathbf{\hat{b}}_{\mathbf{p},y}$> w.r.t. $\mathbf{\hat{b}}_{\mathbf{p},z}$ is defined as follows:
@@
where $\operatorname{sine}(\cdot)$ is a sinusoidal function and $\sigma_{a}$ is a controller constant, all $k$ neighbours are considered to obtain the angular embedding $\mathbf{A}_{\mathbf{p}}$. We update the global geometric embedding $\mathbf{\hat{f}}_{\mathbf{p}}^{gg}$ as an angular-aware attention mechanism:
@@
$\begin{split}\operatorname{Att}(\mathbf{\hat{f}}_{\mathbf{p}}^{g},\mathbf{A}_{\mathbf{p}})=\\ (\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{V}}).\frac{(\mathbf{A}_{\mathbf{p}}\mathbf{W^{A}})(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{Q}})^{\text{T}}+(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{Q}})(\mathbf{\hat{f}}_{\mathbf{p}}^{g}\mathbf{W^{K}})^{\text{T}}}{\sqrt{dim}}.\end{split}$
$\mathbf{W^{A}},\mathbf{W^{Q}},\mathbf{W^{K}},\mathbf{W^{V}}\in\mathbb{R}^{d\times d}$ are the projection matrices of each item and LayerNorm [2] is applied to Eq. 7. Each local point feature is associated with its corresponding global embedding by $\mathbf{\widetilde{f}}_{\mathbf{p}}=\mathbf{f}_{\mathbf{p}}\oplus\mathbf{\hat{f}}_{\mathbf{p}}^{gg},\mathbf{\widetilde{f}}_{\mathbf{p}}\in\mathbb{R}^{N\times 2d}$ to obtain $\mathcal{\widetilde{P}}=\{\mathbf{b}_{\mathbf{p}},\mathbf{\widetilde{f}}_{\mathbf{p}}\}$. The same procedure obtains local and global embedding $\mathcal{\widetilde{Q}}=\{\mathbf{b}_{\mathbf{q}},\mathbf{\widetilde{f}}_{\mathbf{q}}\}$ for the point cloud $\mathcal{\hat{Q}}$.